A body of mass 4kg is on the point of slipping down the plane inclined at 30°. What force parallel parking the plane will just move it up the plane?


1
Expert's answer
2021-02-15T17:02:57-0500
"F=mg\sin{30}\\F=(4)(9.8)\sin{30}=19.6\ N"